Case Study
A travel company has hired a management consulting company to analyze demand in 26 regional markets for one of its major products: a guided tour to a particular country. The consultant uses data to estimate the following equation:
Q = 1,500 = 4P + 5A+ 10I +3PX
where Q = amount of the product demanded
P = price of the product in dollars
A= advertising expenditures in thousands of dollars.
I = income in thousands of dollars
PX = price of some other travel products offered by a competing travel company
Questions:
a. Calculate the amount demanded for this product using the following data:
P = $400
A= $20,000
GM04/July 2014 Page 2/5
I = $15,000
PX = $500 (5)
b. Suppose the competitor reduced the price of its travel products to $400 to match the price of this firm’s product. How much would this firm have to increase its advertising in order to counteract the drop in its competitor’s price? Would it be worth it for them to do so? Explain. (10)
c. What other variables might be important in helping estimate
1
Expert's answer
2014-10-14T12:36:09-0400
Q = 1,500 + 4P + 5A+ 10I +3PX
a. Calculate the amount demanded for this product using the following data:
P = $400, A= $20,000, I = $15,000, PX = $500
Q = 1,500 - 4*400 + 5*20,000 + 10*15,000 + 3*500 = 251,400 units
b. If the competitor reduced the price of its travel products to $400 to match the price of this firm’s product, the total demand for firm's product will decrease by 3*(500 - 400) = 300 units. So, this firm have to increase its advertising by 300/5 = 60 in order to counteract the drop in its competitor’s price.
c. Such other variables as change in tastes, popularity of the product, change in consumer's income might be important in helping estimate the demand.